0 0
Add a comment Improve this question Transcribed image text
Answer #1

since the light is emitted from shere having radius R1 hence this radius remains as it is

the velocity of light i vaccum=c -m/s

but in any medium having refractive index n,the velocity is =c/n -m/s

that is velocity get decrease by 1/n times of vaccum

time taken to cover R2 distance, t=R2/(c/n) seconds

observing from far away the equivalent distance for the observer should be = R1+t*C=R1+nR2

hence equivalent distance =R1+nR2